Minutes of Rosen07 Preparation meeting

+ some post-meeting inputs

March 14, 2007



  • Hall C reconfiguration

           - The main issue in the Hall C reconfiguration is the move and warm-up of G0 magnet.
                --> Smallest angle needed is 10.5o: might need to cut a notch in the target plateform.
                --> G0 warm-up will start as soon as target is cooldown.
           - The priority is to remove G0 target to leave target group work as soon as possible.
           - Install upstream girder.
           - Need people to move green wall blocks.
           - Install small scattering angle beam pipe.
           - SOS hoses replacement and also some in HMS (few hours)
           - HMS magnet cool down (about 2 weeks needed)
           - Tests of the HMS controls.

  • Target

           - The target chamber is here and looks acceptable.
           - Cryotarget cells should be still OK and will be rechecked.
           - The optics target needs some modifications.
           - Dummy Aluminum target: changes in the design. John will look into it.
           - Jan05 solid target ladder still exists but it is no sure if e can re-use it. Eric would rather have a new one (problem with centering with the old one).
           - List of solid targets can be found on Dave Meekins' webpage : Jan05 target setup.
           - Jan05 solid targets still exist and have been remeasured: some have a little difference: New thicknesses.
           - New TSOP with the new target chamber.

  • Beamline

           - Check with Paul Brindza to put exit window in to be able to take 100μA.
           - No valve: might interfere with small angle setup. Downstream beam pipe will be directly connected to the scattering chamber.
           - Raster setup (actually set to "G0 mode")
           - Make sure to check the raster setup after each energy change.

  • HMS Detectors

           - Drift chamber: fast bus not working yet (no power supply)
           - DAQ working. Need to make sure old HV won't be restored.
           - Hodoscopes: gain matching is done. Only 4 channels with problems: will be checked when Hall will be open.
           - For all detectors, need cosmics test to give a real diagnostic.
           - Analyzer: will use the basic Hall C code.

  • SOS Detectors

           - For parasitic, only tracking and triggering are needed.
           - Hamlet and collaborators will work on the SOS when HMS will be ready to take cosmic.

  • Surveys

           - Coordinate directly with Walter so they won't interfere with his work.
           - Beam pipe and upstream girder.
           - Target. Then monitoring of the target position shift as it cooldown.
           - HMS pointing: 4 small angles (including the smallest achievable) and 2 large angles to be surveyed.

  • Website and Shifts

           - First target training session: April 26, 2007.
           - Training for the new HMS controls.
           - Documentation: COO, Radiation budget, ASAD: one set for the three experiments.
           - Shift checklist: add hodoscope efficiency monitoring.
           - Commissionning and production runplans

  • Manpower

           - PDL: Mark Jones
           - Run coordinators needed
           - Graduate Students: Ibrahim Albayrak (Hampton U.), James Johnson (Northwestern U.), Vahe Mamyan (UVA), Georgie Mbianda Njencheu (University of the Witwatersrand)

  • Parasitic tests

           - Do we need to man the shifts during the sHMS test ?
           - Peter Bosted: test of a scintillator array of 64 blocks and a lucite bar in the SOS.
           - Sam Danagoulian: scintillator test in the HMS. Should be OK if put after calorimeters.


    Patricia Solvignon
    Last modified: Thu Mar 22 16:58:48 EDT 2007